OpenStudy (girlmeetsworld):

use inverse operations to solve each equation. explain each step. 19=h/3-8

OpenStudy (girlmeetsworld):

@AloneS

OpenStudy (alexandervonhumboldt2):

add 8 to both sides: 19+8=h/3 -8 +8

OpenStudy (girlmeetsworld):

the video that explained how to do this got me confused so i really need help. ASAP!

OpenStudy (alexandervonhumboldt2):

h/3=27

Mehek (mehek14):

first add 19 + 8 then multiply by 3

OpenStudy (alexandervonhumboldt2):

multply by 3: h=27*3=81

OpenStudy (anonymous):

-95/19=-5 3-8=-5 h=-95

OpenStudy (girlmeetsworld):

where did u get -95 from?

OpenStudy (anonymous):

-5x19=-95 thats how :D

Mehek (mehek14):

where did u get -5 from?

OpenStudy (alexandervonhumboldt2):

you are worng @AskMeLotss

OpenStudy (anonymous):

because it says h/-5=19 so do the opposite -5x19=-95 :)

OpenStudy (anonymous):

3-8=-5

Mehek (mehek14):

hun h/3 not h/-5

OpenStudy (girlmeetsworld):

alex is that the correct answer?

OpenStudy (girlmeetsworld):

what u said?

OpenStudy (anonymous):

oooh I see

OpenStudy (anonymous):

welp...nvm dont listen to what I said before... XD

OpenStudy (girlmeetsworld):

thanks @AlexandervonHumboldt2

OpenStudy (girlmeetsworld):

and @Mehek14

OpenStudy (anonymous):

I dont know if you need inverse operations but the equations can be solved with the following steps.. Eq 1: 19 = h/3 -8 Add 8 to both sides of equation 19 + 8 = h/3 or h/3 = 27 Multiply both sides by 3 h = 27*3 = 81

OpenStudy (girlmeetsworld):

thanks. :)

OpenStudy (anonymous):

yes sorry, I got confused for a sec, but this is how it goes :)

OpenStudy (girlmeetsworld):

lol its ok. :)

OpenStudy (anonymous):

I know...for a moment it was like, what is this?! XD but yep I see now...thanks @Mehek14

Mehek (mehek14):

yw :)

OpenStudy (anonymous):

lol k

OpenStudy (anonymous):

can I at least get a medal for trying...i solved it now

OpenStudy (anonymous):

:D

OpenStudy (girlmeetsworld):

lol i gave the medal to @AlexandervonHumboldt2. sorry next time though. ok? :)

OpenStudy (girlmeetsworld):

oh what the heck

OpenStudy (girlmeetsworld):

there u go. :)

OpenStudy (girlmeetsworld):

@AskMeLotss

OpenStudy (anonymous):

hello? oh, thanks for the medal XD

OpenStudy (girlmeetsworld):

np

OpenStudy (anonymous):

sorry, i was helping others XD

OpenStudy (girlmeetsworld):

hey ... i need more help

OpenStudy (girlmeetsworld):

what do u do with the 16?

OpenStudy (girlmeetsworld):

@AskMeLotss

OpenStudy (anonymous):

im here to help

OpenStudy (anonymous):

h=81

OpenStudy (girlmeetsworld):

@Mehek14

OpenStudy (anonymous):

it does... h=81

OpenStudy (girlmeetsworld):

if u do 19+8=h/3-8+8 then what do u do with the 16?

OpenStudy (anonymous):

16? what are you talking about?

OpenStudy (girlmeetsworld):

@Mehek14

OpenStudy (girlmeetsworld):

8+8

OpenStudy (girlmeetsworld):

im confused by alex's comment

OpenStudy (girlmeetsworld):

@Mehek14 please explain everything to me.

Mehek (mehek14):

19+8=h/3-8+8 ?

OpenStudy (girlmeetsworld):

thats what alex said to do but ... im not sure thats right.

Mehek (mehek14):

yes so 19 + 8 = 27 27 = h/3 multiply 27 * 3 = ?

OpenStudy (anonymous):

@Mehek14 what do you think he's talking about...its not making sense...

OpenStudy (girlmeetsworld):

im a girl lol.

OpenStudy (anonymous):

I meant she XD

OpenStudy (girlmeetsworld):

yes so 19 + 8 = 27 27 = h/3 multiply 27 * 3 = ? mehek that was ur response and im still a little bit confused.

OpenStudy (girlmeetsworld):

can u explain it a little please mehek

OpenStudy (girlmeetsworld):

@Mehek14

Mehek (mehek14):

when you want to move -8 to the other side, you do the opposite which is +8 -8 + 8 = 0 and 19+8 = 27 so you have 27 = h/3 you want 3 on the other side so you do the opposite of /3 which is *3 /3 * 3 = 0 27 * 3 = 81

OpenStudy (girlmeetsworld):

oh ... i fell really dumb now. :/ thanks

Mehek (mehek14):

ur not dumb..you just need it to be explained for you :) and yw

OpenStudy (girlmeetsworld):

:)